OpenStudy (anonymous):

Solve for x 1-tan^2x = sin^2x + 1/sec^2x

OpenStudy (anonymous):

x=0 or 180 or 360

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

tan(x)=sin(x)/cos(x) , sec(x)=1/cos(x)

OpenStudy (anonymous):

\[1-\tan^2x = \sin^2x + \frac{ 1 }{ \sec^2x } \] \[1-\frac{ sin^2x }{ \cos^2x } = \sin^2x + cos^2x \] \[1-\frac{ sin^2x }{ \cos^2x } =1 \] \[\frac{ sin^2x }{ \cos^2x } =0 \] \[\sin^2x= 0 \]

OpenStudy (anonymous):

\[x=n \pi \] n=0,1,2,.....